# supplementary

### v1

@@ V Limitations
The policy learned with RWM surpasses existing MBRL methods in both robustness and generalization. However, it still falls short of the performance achieved by well-tuned model-free RL methods trained on high-fidelity simulators. Model-free RL, being a more mature and extensively optimized paradigm, excels in settings where unlimited interaction with near-perfect simulators is possible. In contrast, the strengths of MBRL are more pronounced in scenarios where accurate or efficient simulation is infeasible, making it an indispensable tool for enabling intelligent agents to eventually learn and adapt in complex, real-world environments.
@@
In this work, the world model is pre-trained using simulation data prior to policy optimization, reducing instability during training. However, training from scratch remains challenging as policies can exploit model inaccuracies during exploration, leading to inefficiency and instability. In addition, the need for additional interaction with the environment to fine-tune the world model highlights areas for further refinement. Nevertheless, enabling safe and effective online learning directly on hardware remains challenging. Current training in simulation avoids potential hardware damage, but incorporating safety constraints and robust uncertainty estimates will be critical for deploying RWM and MBPO-PPO in real-world, lifelong learning scenarios. These limitations underscore the trade-offs inherent in MBRL frameworks, balancing data efficiency, safety, and performance while addressing the complexities of real-world robotic systems.
VI Conclusion
@@
In this work, we present RWM, a robust and scalable framework for learning world models tailored to complex robotic tasks. Leveraging a dual-autoregressive mechanism, RWM effectively addresses key challenges such as compounding errors, partial observability, and stochastic dynamics. By incorporating historical context and self-supervised training over long prediction horizons, RWM achieves superior accuracy and robustness without relying on domain-specific inductive biases, enabling generalization across diverse tasks.
Through extensive experiments, we demonstrate that RWM consistently outperforms state-of-the-art approaches like RSSM and transformer-based architectures in autoregressive prediction accuracy across diverse robotic environments. Its ability to maintain stability under noise and dynamically complex scenarios further underscores its robustness and applicability to real-world tasks.
Building on RWM, we propose MBPO-PPO, a policy optimization framework that leverages long world model rollout fidelity. Policies trained using MBPO-PPO demonstrate superior performance in simulation and transfer seamlessly to hardware, as evidenced by zero-shot deployment on the ANYmal D robot. Unlike baseline methods such as SHAC and Dreamer, which struggle with unstable dynamics and compounding errors, MBPO-PPO reliably optimizes policies that generalize effectively to physical systems.
This work advances the field of model-based reinforcement learning by providing a generalizable, efficient, and scalable framework for learning and deploying world models. The results highlight RWM ’s potential to enable adaptive, robust, and high-performing robotic systems, setting a foundation for broader adoption of model-based approaches in real-world applications.
Appendix A Appendix
A-A Task Representation
A-A1 Observation and action spaces
@@
The observation space for the ANYmal world model is composed of base linear and angular velocities $v$, $\omega$ in the robot frame, measurement of the gravity vector in the robot frame $g$, joint positions $q$, velocities $\dot{q}$ and torques $\tau$ as in Table S1.
@@
Entry Symbol Dimensions
base linear velocity $v$ 0:3
base angular velocity $\omega$ 3:6
projected gravity $g$ 6:9
joint positions $q$ 9:21
joint velocities $\dot{q}$ 21:33
joint torques $\tau$ 33:45
@@
Entry Symbol Dimensions
knee contact -- 0:4
foot contact -- 4:8
@@
Entry Symbol Dimensions
joint position targets $q^{*}$ 0:12
@@
Entry Symbol Dimensions
base linear velocity $v$ 0:3
base angular velocity $\omega$ 3:6
projected gravity $g$ 6:9
velocity command $c$ 9:12
joint positions $q$ 12:24
joint velocities $\dot{q}$ 24:36
A-A2 Reward functions
@@
Symbol $w_{v_{xy}}$ $w_{\omega_{z}}$ $w_{v_{z}}$ $w_{\omega_{xy}}$ $w_{q_{\tau}}$
Value $1.0$ $0.5$ $-2.0$ $-0.05$ $-2.5e^{-5}$
Symbol $w_{\ddot{q}}$ $w_{\dot{a}}$ $w_{f_{a}}$ $w_{c}$ $w_{g}$
Value $-2.5e^{-7}$ $-0.01$ $0.5$ $-1.0$ $-5.0$
@@
where $c_{u}$ denotes the counts of the undesired knee contacts.
@@
A-B Network Architecture
A-B1 RWM
The robotic world model consists of a GRU base and MLP heads predicting the mean and standard deviation of the next observation and privileged information such as contacts, as detailed in Table S6.
@@ A-B2 Baselines
A-B3 MBPO-PPO
The network architectures of the policy and the value function used in MBPO-PPO are detailed in Table S8.
@@ A-C1 RWM
max iterations -- $2500$
learning rate -- $1e^{-4}$
weight decay -- $1e^{-5}$
batch size -- $1024$
@@
approximate training hours -- $1$
number of seeds -- $5$
A-C2 MBPO-PPO
@@
imagination environments -- $4096$
imagination steps per iteration -- $100$
@@
max iterations -- $2500$
learning rate -- $0.001$
weight decay -- $0.0$
learning epochs -- $5$
mini-batches -- $4$
KL divergence target -- $0.01$
@@
entropy coefficient -- $0.005$
number of seeds -- $5$

### v4

@@ 5 Limitations
The policy learned with RWM and MBPO-PPO surpasses existing MBRL methods in both robustness and generalization. However, it still falls short of the performance achieved by well-tuned model-free RL methods trained on high-fidelity simulators. Model-free RL, being a more mature and extensively optimized paradigm, excels in settings where unlimited interaction with near-perfect simulators is possible. In contrast, the strengths of MBRL are more pronounced in scenarios where accurate or efficient simulation is infeasible, making it an indispensable tool for enabling intelligent agents to eventually learn and adapt in complex, real-world environments. To clarify the computational and performance aspects, we provide a comparison against a PPO-based method with a high-fidelity simulator in Table 1.
Table 1: Comparison with model-free method
Method RWM pretraining MBPO-PPO PPO
state transitions 6M - 250M
total training time 50 min 5 min 10 min
step inference time - 1 ms 1 ms
real tracking reward - $0.90\pm 0.04$ $0.90\pm 0.03$
@@
In this work, the world model is pre-trained using simulation data prior to policy optimization, reducing instability during training (see Sec. A.4.3). However, training from scratch remains challenging as policies can exploit model inaccuracies during exploration, leading to inefficiency and instability. In addition, the need for additional interaction with the environment to fine-tune the world model highlights areas for further refinement. Nevertheless, enabling safe and effective online learning directly on hardware remains challenging (see Sec. A.4.4). Current training in simulation avoids potential hardware damage, but incorporating safety constraints and robust uncertainty estimates will be critical for deploying RWM and MBPO-PPO in real-world, lifelong learning scenarios. These limitations underscore the trade-offs inherent in MBRL frameworks, balancing data efficiency, safety, and performance while addressing the complexities of real-world robotic systems.
6 Conclusion
@@
In this work, we present RWM, a robust and scalable framework for learning world models tailored to complex robotic tasks. Leveraging a dual-autoregressive mechanism, RWM effectively addresses key challenges such as compounding errors, partial observability, and stochastic dynamics. By incorporating historical context and self-supervised training over long prediction horizons, RWM achieves superior accuracy and robustness without relying on domain-specific inductive biases, enabling generalization across diverse tasks. Through extensive experiments, we demonstrate that RWM consistently outperforms state-of-the-art approaches like RSSM and transformer-based architectures in autoregressive prediction accuracy across diverse robotic environments. Building on RWM, we propose MBPO-PPO, a policy optimization framework that leverages long world model rollout fidelity. Policies trained using MBPO-PPO demonstrate superior performance in simulation and transfer seamlessly to hardware, as evidenced by zero-shot deployment on the ANYmal D and Unitree G1 robots. This work advances the field of model-based reinforcement learning by providing a generalizable, efficient, and scalable framework for learning and deploying world models. The results highlight RWM ’s potential to enable adaptive, robust, and high-performing robotic systems, setting a foundation for broader adoption of model-based approaches in real-world applications.
Acknowledgments and Disclosure of Funding
This research was supported by the ETH AI Center.
Appendix A Technical Appendices and Supplementary Material
A.1 Task Representation
A.1.1 Observation and action spaces
@@
The observation space for the ANYmal D and Unitree G1 world model is composed of base linear and angular velocities $v$, $\omega$ in the robot frame, measurement of the gravity vector in the robot frame $g$, joint positions $q$, velocities $\dot{q}$ and torques $\tau$ as in Table S2.
@@
Entry Symbol Dimensions Entry Symbol Dimensions
ANYmal D Unitree G1
base linear velocity $v$ 0:3 base linear velocity $v$ 0:3
base angular velocity $\omega$ 3:6 base angular velocity $\omega$ 3:6
projected gravity $g$ 6:9 projected gravity $g$ 6:9
joint positions $q$ 9:21 joint positions $q$ 9:38
joint velocities $\dot{q}$ 21:33 joint velocities $\dot{q}$ 38:67
joint torques $\tau$ 33:45 joint torques $\tau$ 67:96
@@
Entry Symbol Dimensions Entry Symbol Dimensions
ANYmal D Unitree G1
knee contact - 0:4 body contact - 0:26
foot contact - 4:8 foot height - 26:28
foot velocity - 28:30
@@
Entry Symbol Dimensions Entry Symbol Dimensions
ANYmal D Unitree G1
joint position targets $q^{*}$ 0:12 joint position targets $q^{*}$ 0:29
@@
Entry Symbol Dimensions Entry Symbol Dimensions
ANYmal D Unitree G1
base linear velocity $v$ 0:3 base linear velocity $v$ 0:3
base angular velocity $\omega$ 3:6 base angular velocity $\omega$ 3:6
projected gravity $g$ 6:9 projected gravity $g$ 6:9
velocity command $c$ 9:12 velocity command $c$ 9:12
joint positions $q$ 12:24 joint positions $q$ 12:41
joint velocities $\dot{q}$ 24:36 joint velocities $\dot{q}$ 41:70
A.1.2 Reward functions
@@
Symbol Value Symbol Value Symbol Value Symbol Value
ANYmal D Unitree G1
$w_{v_{xy}}$ $1.0$ $w_{\omega_{z}}$ $0.5$ $w_{v_{xy}}$ $1.0$ $w_{\omega_{z}}$ $0.5$
$w_{v_{z}}$ 2.0-2.0 $w_{\omega_{xy}}$ 0.05-0.05 $w_{v_{z}}$ 2.0-2.0 $w_{\omega_{xy}}$ 0.05-0.05
$w_{q_{\tau}}$ $2.5e^{-5}$ $w_{\ddot{q}}$ $2.5e^{-7}$ $w_{q_{\tau}}$ $2.5e^{-5}$ $w_{\ddot{q}}$ $2.5e^{-7}$
$w_{\dot{a}}$ 0.01-0.01 $w_{f_{a}}$ $0.5$ $w_{\dot{a}}$ 0.05-0.05 $w_{f_{a}}$ $0.0$
$w_{c}$ 1.0-1.0 $w_{g}$ 5.0-5.0 $w_{c}$ 1.0-1.0 $w_{g}$ 5.0-5.0
$w_{f_{c}}$ $0.0$ $w_{q_{d}}$ $0.0$ $w_{f_{c}}$ $1.0$ $w_{q_{d}}$ 1.0-1.0
@@
where $c_{u}$ denotes the counts of the undesired contacts.
@@
Foot clearance
$r_{f_{c}}=w_{f_{c}}h_{f_{c}},$
where $h_{f_{c}}$ denotes the clearance height of the swing feet.
Joint deviation
$r_{q_{d}}=w_{q_{d}}\left\|q-q_{0}\right\|_{1},$
where $q_{0}$ denotes the default joint position.
A.2 Network Architecture
A.2.1 RWM
The robotic world model consists of a GRU base and MLP heads predicting the mean and standard deviation of the next observation and privileged information such as contacts, as detailed in Table S7. The training scheme is visualized in Fig. S6.
Figure S6: Dual-autoregressive mechanism employed in RWM. Inner autoregression updates GRU hidden states after each historical step within the context horizon, while outer autoregression feeds predicted observations from the forecast horizon back into the network. The dashed arrows denote the sequential autoregressive prediction steps, highlighting robustness to long-term dependencies and transitions.
@@ A.2.2 Baselines
A.2.3 MBPO-PPO
The network architectures of the policy and the value function used in MBPO-PPO are detailed in Table S9. The training scheme is visualized in Fig. S7.
Figure S7: Model-Based Policy Optimization with learned world models. The framework combines real environment interactions with simulated rollouts for efficient policy optimization. Observation and action pairs from the environment are stored in a replay buffer and used to train the autoregressive world model. Imagination rollouts using the learned model predict future states over a horizon of $T$, providing trajectories for policy updates through reinforcement learning algorithms.
@@ A.3.1 RWM
max iterations - $2500$
learning rate - $1e^{-4}$
weight decay - $1e^{-5}$
batch size - $1024$
@@
approximate training hours - 11
number of seeds - 55
A.3.2 MBPO-PPO
@@
imagination environments - $4096$
imagination steps per iteration - $100$
@@
max iterations - $2500$
learning rate - $0.001$
weight decay - $0.0$
learning epochs - 55
mini-batches - 44
KL divergence target - $0.01$
@@
entropy coefficient - $0.005$
number of seeds - 55
A.4 Additional Experiments and Discussions
A.4.1 Dual-autoregressive Mechanism
The heatmap on the left in Fig. S8 shows the relative autoregressive prediction error $e$ under different combinations of $M$ and $N$. Models trained with a longer history horizon $M$ consistently exhibit lower prediction errors, demonstrating the importance of providing sufficient historical context to capture the underlying dynamics. However, the influence of $M$ plateaus beyond a certain point, indicating diminishing returns for very large history horizons. Forecast horizon $N$, on the other hand, plays a decisive role in improving long-term prediction accuracy. Increasing $N$ during training leads to better performance in autoregressive rollouts, as it encourages the model to learn representations robust to compounding errors over extended prediction horizons. This improvement comes at the cost of increased training time, as shown in the heatmap on the right. Larger $N$ values require sequential computation during training due to the autoregressive nature of the process, significantly lengthening the training duration.
Figure S8: Ablation study on the history horizon $M$ and forecast horizon NN in RWM. The heatmap on the left shows the relative autoregressive prediction error, with darker colors indicating higher errors. Models trained with larger history horizons $M$ exhibit lower errors, although the improvements plateau beyond a certain point. Forecast horizon NN has a significant impact, with longer horizons leading to better long-term prediction accuracy due to exposure to extended rollouts during training. The heatmap on the right illustrates training time, with darker colors representing longer durations. Increasing NN significantly raises training time due to sequential computation, while shorter horizons (e.g., $N=1$, teacher-forcing) enable faster training but result in poor prediction accuracy.
A.4.2 Visualization of Imagination Rollouts
The imagination rollouts across various robotic environments compared with the ground-truth simulation is visualized in Fig. S9.
Figure S9: Autoregressive imagination of RWM and ground-truth simulation across diverse robotic systems. For each environment, the top row showcases the RWM autoregressively predicting future trajectories in imagination. The second row visualizes the ground truth evolution in simulation. The visualized coordinate and arrow markers denote the predicted and measured end-effector pose and base velocity, respectively.
A.4.3 Collision Handling and Model Pretraining
In both phases of the pretraining and online fine-tuning of RWM, we terminate rollouts and reset the environment when ground contact by the base is detected, signaling a failure. We explicitly train RWM to predict such terminations in its privileged information prediction head. This enables the world model to learn transitions leading to unsafe situations. During policy optimization, MBPO-PPO treats these termination predictions as episode-ending events in imagination rollouts, affecting PPO’s return computation and state values.
RWM is pretrained with simulation data induced by policies trained for similar tasks under varied dynamics. The policy is learned from scratch purely in imagination, with RWM fine-tuned using a single-environment online dataset. Pretraining is essential for two key reasons. First, the online dataset is extremely limited, as it is generated by only a single environment, akin to real-world constraints. Training the world model entirely from scratch on such data would lead to severe overfitting and long training times. Second, an immature policy would frequently cause the robot to fall, generating transitions with limited value. In cases of significant failure or domain shift, training the world model solely on these data would result in chaotic imagined rollouts, which in turn would produce poor policy updates. Pretraining stabilizes training and serves as a robust initialization for online fine-tuning, particularly in environments with challenging dynamics.
Importantly, RWM pretraining does not require data from optimal policies. Figure 3 demonstrate that RWM remains robust to domain shifts and injected noise. As an alternative, we warm up the model using data from a suboptimal policy, which significantly stabilizes training. Notably, this pretraining is only necessary for locomotion tasks due to the discontinuous dynamics and environment terminations. Our manipulation experiments do not require such pretraining.
A.4.4 Challenges in Real-World Online Learning
We acknowledge that the advantages of our approach would be further demonstrated by performing the policy training phase directly on real hardware. While this is a key long-term objective, several challenges currently prevent real-world deployment.
During online learning, the policy often exploits minor world model errors, leading to overly optimistic behaviors that result in collisions. In simulation, these failures serve as corrective signals, but in real hardware, they pose a risk to the robot. Our experiments show that such failures occur more than 20 times on average during online learning, which would be detrimental to real-world systems. Even if hardware collisions were acceptable, fully automating online learning would require a recovery policy capable of resetting the robot to an initial state—a particularly challenging requirement for large platforms like ANYmal D or Unitree G1. Additionally, privileged information used to fine-tune RWM (e.g., contact forces) must be either measured or estimated using onboard sensors, which may not always be available. To mitigate error exploitation, uncertainty-aware world models could be explored, but integrating such models into RWM would require additional architectural modifications. Due to these challenges, we approximate real-world constraints by using only a single simulation environment with domain shifts from pretraining environments. This setup reduces engineering effort while proving the feasibility of our approach. Our ongoing work specifically addresses these issues .
A.5 Ethics and Societal Impacts
This work does not involve human subjects or sensitive data. All experiments are conducted in simulation or on dedicated robotic hardware operated by the authors, with no use of third-party datasets. The research complies with the Code of Ethics of the venue. The proposed framework provides a robust and scalable method for learning world models tailored to complex robotic tasks. This can benefit domains such as healthcare, disaster response, and logistics, and reduce environmental and hardware costs associated with physical experimentation. Potential risks include misuse of the method in surveillance or autonomous enforcement systems, and the acceleration of automation in labor-sensitive sectors. While such uses are not intended or explored in this work, the authors acknowledge the dual-use potential of generalizable control methods. To mitigate safety risks, policy training occurs entirely in simulation, and deployment is limited to policies validated under domain shifts. Failure events are explicitly modeled and used to terminate unsafe rollouts. Online learning on hardware is deferred due to safety concerns and the absence of reliable recovery strategies. Future work will explore uncertainty-aware models and safer online adaptation.